# Learning About Social Context from Smartphone Data: Generalization Across Countries and Daily Life Moments

### 29 Jan 2024 v4

@@ 5.2. RQ2: Social Context Inference Without Personalization
Figure 9. Comparison of random forest model performance for population-level (leave-k-out) and hybrid (personalized) models. The performance metric shown is F1 score.
@@ 5.4. RQ4: Generalization of Models to Unseen Countries
Testing UK Denmark Italy Paraguay Mongolia
Training
UK 67.1 (4.0) / 61.5 (3.2) 54.2 (10.0) / 50.1 (8.1) 60.3 (3.1) / 56.7 (1.8) 61.8 (8.9) / 55.6 (6.9) 53.8 (3.9) / 42.5 (5.9)
Denmark 51.9 (4.2) / 47.0 (3.8) 67.8 (9.2) / 58.4 (9.6) 55.4 (3.3) / 49.0 (2.7) 56.0 (7.8) / 51.7 (5.8) 52.2 (6.8) / 50.7 (3.9)
Italy 61.9 (4.3) / 55.8 (1.8) 60.4 (10.8) / 54.5 (11.8) 66.9 (2.6) / 61.5 (2.0) 54.2 (6.6) / 50.8 (7.0) 57.6 (5.1) / 42.0 (8.5)
Paraguay 58.0 (7.1) / 51.8 (7.5) 50.1 (11.2) / 43.8 (9.4) 56.1 (3.3) / 53.1 (2.1) 75.4 (14.4) / 68.5 (13.0) 57.4 (5.1) / 48.8 (4.5)
Mongolia 55.5 (6.8) / 31.0 (11.1) 54.4 (13.9) / 45.4 (8.4) 52.8 (3.7) / 34.8 (4.5) 62.8 (9.8) / 47.5 (13.7) 75.5 (5.4) / 68.7 (6.1)
@@ 6.1. Summary of Results
  • RQ1: We conducted a comprehensive analysis of a smartphone sensing dataset collected from 483 young adults across five countries. Our analysis identified specific individual features, such as types and quantities of app usage (including tools, communication, productivity, and social), activity types (like being in a vehicle, on foot, walking), location altitude, Bluetooth and WiFi usage, proximity, and screen episodes, as top indicators for distinguishing between being alone and not alone. Moreover, we found that the features significant in differentiating social contexts varied across countries, emphasizing the diversity in behaviors and contextual features associated with various social settings.
@@
  • RQ3: The utilization of hybrid models resulted in the multi-country approach yielding an AUC of 95.3%. Country-specific models performed similarly, with their differences in AUC being in the range of 3%. The results indicate that a generic multi-country model may generally suffice for social context inference, regardless of personalization level. Country-specific models might offer slight improvements in certain countries (i.e., UK, Denmark, and Mongolia). Notably, this deviates from previous research utilizing smartphone sensor data for health-related inferences, where country-specific models were recommended to achieve better performance, even when personalized. This contrast highlights the differing effectiveness of model types across various domains and objectives.
@@ 8. Conclusion
Appendix A Features Extracted from Smartphone Data
Table 6. Summary tables of sensors and extracted sensor features by authors of the paper that conducted feature processing. The sensor features are computed by aggregating ten minutes of sensor data.
Sensor # of Features and Description
Features
Location (GPS)
3
Radius of gyration, the sum of distance covered, altitude
Bluetooth (low energy)
5
Number of devices connected, mean rssi (received signal strength indicator), max rssi, min rssi, std rssi
Bluetooth (Normal)
5
Number of devices connected, mean rssi (received signal strength indicator), max rssi, min rssi, std rssi
Wifi
6
Number of devices connected to the device wifi hotspot, if the device is connected to a wifi, wifi mean rssi (received signal strength indicator), max rssi, min rssi, std rssi
Cellular gsm (2G)
4
Strength of the mobile signal as defined by mean, max, min, std
Cellular wcdma (3G)
4
Strength of the mobile signal as defined by mean, max, min, std
Cellular LTE (4G)
4
Strength of the mobile signal as defined by mean, max, min, std
Notifications
4
notifications posted, notifications removed, notifications posted without duplicates, notifications removed without duplicates
Proximity
4
Measures of the proximity sensor as mean, max, min, std
Activity
8
Activities as classified from the accelerometer data by the Google Activity Recognition API (Developers, 2021); activity still, activity tilting, activity invehicle, activity onbicycle, activity onfoot, activity walking, activity running, activity unknown
Steps
2
step counted, steps detected
Touch
1
Number of touch events
Screen
7
User presence time, screen number of episodes, screen time total, screen time per episode, screen time max episode, screen time min episode, screen time std episode
Apps Usage
48
Time spent using app categories. App categories as classified by the google playstore (Santani et al., 2018). E.g.: Action apps, dating apps, music apps, puzzle apps, etc.
Time
2
Hour of the day, day of the week

### 01 Mar 2024 v5

@@ 5.2. RQ2: Social Context Inference Without Personalization
The table displays the performance of various machine learning models, including their AUC and F1 scores with standard deviations, in predicting social context across different countries. It highlights that random forest classifiers generally achieved the best performance in terms of both AUC and F1 scores across many countries.
@@
Figure 9. Comparison of random forest model performance for population-level (leave-k-out) and hybrid (personalized) models. The performance metric shown is the F1 score.
This figure compares the performance, in terms of AUC, of random forest models with and without feature selection. It shows that multi-country models perform similarly to country-specific models.
This figure similarly compares the performance of random forest models with and without feature selection, but in terms of the F1 score. It shows that multi-country models are on par with country-specific models.
This figure compares the performance, measured by AUC, of random forest models for population-level (leave-k-out) and hybrid (personalized) models. It demonstrates that hybrid models achieve high AUCs above 90% in both country-specific and multi-country setups.
This figure compares the performance, measured by F1 score, of random forest models for population-level (leave-k-out) and hybrid (personalized) models. It shows that hybrid models achieve high F1 scores above 85% in both country-specific and multi-country setups.
@@
Figure 9. Comparison of random forest model performance for population-level (leave-k-out) and hybrid (personalized) models. The performance metric shown is the F1 score.
@@ 5.4. RQ4: Generalization of Models to Unseen Countries
[…]
This table shows the performance of country-agnostic models in predicting social context in different countries, using AUC and F1 scores with their standard deviations. The diagonal values, which are grayed out, indicate the performance obtained from the country-specific approach with feature selection, as shown in Figure 6. Results show that social context inference models do not generalize well to other countries. Testing UK Denmark Italy Paraguay Mongolia Training UK 67.1 (4.0) / 61.5 (3.2) 54.2 (10.0) / 50.1 (8.1) 60.3 (3.1) / 56.7 (1.8) 61.8 (8.9) / 55.6 (6.9) 53.8 (3.9) / 42.5 (5.9) Denmark 51.9 (4.2) / 47.0 (3.8) 67.8 (9.2) / 58.4 (9.6) 55.4 (3.3) / 49.0 (2.7) 56.0 (7.8) / 51.7 (5.8) 52.2 (6.8) / 50.7 (3.9) Italy 61.9 (4.3) / 55.8 (1.8) 60.4 (10.8) / 54.5 (11.8) 66.9 (2.6) / 61.5 (2.0) 54.2 (6.6) / 50.8 (7.0) 57.6 (5.1) / 42.0 (8.5) Paraguay 58.0 (7.1) / 51.8 (7.5) 50.1 (11.2) / 43.8 (9.4) 56.1 (3.3) / 53.1 (2.1) 75.4 (14.4) / 68.5 (13.0) 57.4 (5.1) / 48.8 (4.5) Mongolia 55.5 (6.8) / 31.0 (11.1) 54.4 (13.9) / 45.4 (8.4) 52.8 (3.7) / 34.8 (4.5) 62.8 (9.8) / 47.5 (13.7) 75.5 (5.4) / 68.7 (6.1)
@@
This figure illustrates the performance of country-agnostic models using random forest classifiers, with AUC and F1 scores for different testing countries. It highlights that these models consistently score lower in AUC compared to country-specific models, though the differences are generally minor.
@@ 6.1. Summary of Results
  • RQ1: We conducted a comprehensive analysis of a smartphone sensing dataset collected from 483 young adults across five countries. Our analysis identified specific individual features, such as types and quantities of app usage (including tools, communication, productivity, and social), activity types (like being in a vehicle, on foot, walking), location, Bluetooth and WiFi usage, proximity, and screen episodes, as top indicators for distinguishing between being alone and not alone. Moreover, we found that the features significant in differentiating social contexts varied across countries, emphasizing the diversity in behaviors and contextual features associated with various social settings.
@@
  • RQ3: The utilization of hybrid models resulted in the multi-country approach yielding an AUC of 95.3%. Country-specific models performed similarly, with their differences in AUC being in the range of 3%. The results indicate that a generic multi-country model may generally suffice for social context inference, regardless of personalization level. Country-specific models might offer slight improvements in certain countries (i.e., UK, Denmark, and Mongolia). Notably, this deviates from previous research that use smartphone sensor data for health-related inferences, where country-specific models were recommended to achieve better performance, even when personalized. This contrast highlights the differing effectiveness of model types across various domains and objectives.
@@ 8. Conclusion
Acknowledgements.
This work was funded by the European Union’s Horizon 2020 WeNet project, under grant agreement 823783. We deeply thank all the volunteers across the world for their participation in the study.